OpenStudy (anonymous):

OK I really need help on this problem, im stumped! Solve the system of linear equations by graphing 2x-y = 0 3x - y = -1

OpenStudy (anonymous):

Hi you have to plot both lines. so rearrange both into the form y=...

OpenStudy (anonymous):

so y=2x - 0 y= 3x- -1? Im sorry im really slow in math..LOL

OpenStudy (anonymous):

so yh y=2x-0 so y=2x. y=3x- -1=3x +1. now u have to plot these two lines. you ever done this?

OpenStudy (anonymous):

No i havent, LOL so I add 3x + 1= 4x and 2x+0 = 2x? so the plots are 4, 2? if so which one so i put on the x and y?

OpenStudy (anonymous):

ok lets start with y=2x. if I was you I would pick 5 values of x to sub in. so say x=-2, y= -4. x=-1 , y=-2... etc

OpenStudy (anonymous):